OpenStudy (anonymous):

Find value of: cot^-1 (4) (inverse of cotangent) Why do we have to use cosine to evaluate this? Can someone explain clearly please? Why not just put the inverse of cot into the calculator?

OpenStudy (ash2326):

@MathCurious are you here?

OpenStudy (anonymous):

i am here? Could you help me?

OpenStudy (ash2326):

We need to find the angle. Does your text mentions to use cosine?

OpenStudy (anonymous):

What I understand is I have to use the given values from the Cotangent to find the hyp. then use the Cosine function. The only mention of this the text makes is because they have the same domain [0,pi) but I don't see why I can't just use cotangent in my calculator? I have to evaluate it using Cosine.

OpenStudy (ash2326):

Even if we find cosine, we can't find angle without a trigonometric table or calculator. Do you have a trigonometric table for cosine or sin

OpenStudy (anonymous):

According to the cot the X and Y are given. So if cot (theta) = 4 X=4 Y=1

OpenStudy (ash2326):

it's becaus cot, sec and cosec are not in calculators. First we need to find cosine and then we can find using calculator

OpenStudy (anonymous):

but isn't cot just 1 / tan ? calculators have tangent.

OpenStudy (ash2326):

we could do that too :)

OpenStudy (anonymous):

If we use cosine it is (\[4\sqrt{17}\div17\] This isn't the same as the inverse cotangent of 4.So we cannot use 1 / tan

OpenStudy (ash2326):

no it'd be \[\cos^{-1} \frac {4}{\sqrt {17}}\]

OpenStudy (ash2326):

we can use tan, no problem with that

OpenStudy (anonymous):

according to my book and homework it is cos-1 (4 sqrt(17) / 17 ) You have to rationalize the denominator. I have the correct answer, i am just not clear on why it is this way. if you took 1/tan-1(4) you'd ge ta totally different value. hmm..

OpenStudy (ash2326):

Google these two "arccos (4 sqrt 17 /17) in degrees" "arctan ( 1/4) in degrees" Both will give you the same result

OpenStudy (ash2326):

arc cos= inverse cosine

OpenStudy (anonymous):

In google I got different numbers, unfortunately.The question calls for radians

OpenStudy (anonymous):

Ok I added the "in degrees' and got the same for degrees. I must be doing something wrong on my calculator? haha. because 1 / tan-1(4) does not yield the same results..

OpenStudy (ash2326):

yes, it's tan-1 (1/4)

OpenStudy (ash2326):

\[\cot^{-1} (4)\ne \frac{1}{\tan^{-1} 4}\] \[\cot^{-1} (4)= \tan^{-1} \frac 1{4}\]

OpenStudy (anonymous):

Oooh. How did you come up with this? if tan = y/x then... 1 / y/x = x/y. So why doesn't cot = 1 / tan ??? I am so confused, but it's probablyu so simple!

OpenStudy (ash2326):

Just a moment, I'll explain you

OpenStudy (ash2326):

let "a" be the angle \[\tan a=\frac {x}{y}\] \[\cot a =\frac {y}{x}\] hence \[a=\cot^{-1} \frac {y}{x}\] from the first expression \[a=\tan^{-1}\frac{x}{y}\]

OpenStudy (anonymous):

Okay, it's so simple! Thank you so much for your help. Medal!

OpenStudy (ash2326):

Welcome :D